OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). Answer y = 3x - 5 y = 3x - 2 y = 3x + 1 y = 3x + 3

OpenStudy (anonymous):

y = 3x - 5

OpenStudy (anonymous):

Use the slope-intercept formula: \[y-y _{1}=m(x-x_{1})\]\[y-(-2)=3(x-1)\]\[y+2=3x-3\]\[y=3x-5\]

OpenStudy (anonymous):

Thanks!